OpenStudy (anonymous):

-[3x+(7x+4)]=9-(7x+9)

OpenStudy (anonymous):

Distibute the - on the right hand side. you will get: 9-7x-9. This gives you(on the right hand side) : 7x. Then on the left hand side distribute the negative side to all terms in the brackets, resluting in : -3x-7x-4. Giving you on the left hand side: -10x-4. The equation shoud look like this: -10x-4=7x. Now add 10x to both sides and you will get: -4=17x. Finally divide both sides by 17, giving you x=-4/17

OpenStudy (anonymous):

wait a second, the equation will look like this: -10x-4=-7x. I am sorry i missed a negative sign

OpenStudy (anonymous):

then add 10x to both sides giving you, -4=3x. Finally divide by 3 (both side) and get x=-4/3

jhonyy9 (jhonyy9):

-3x-7x-4 =9-7x-9 -10x-4=-7x -4=3x --- x= -4/3